Ver confused with this one Solve sytem of linear equation express in ordered pai (x,y) if possible use substitution or addition method
x- 5/3y=7
4/7x + y=-13/7
thanks, please show me steps
dnn

Answer by stanbon(48564)

Solve sytem of linear equation express in ordered pai (x,y) if possible use substitution or addition method
1st: x- 5/3y=7
2nd: 4/7x + y=-13/7
Rewrite each without the fractions:
3rd: 3x-5y=21
4th: 4x+7y=-13
Multiply 3rd by 4 and 4th by 3 to get:
5th: 12x-20y=84
6th: 12x+21y=-39
Subtract 5th from 6th to get:
41y=--123
y=-3
Substitute into 3rd to solve for x, as follows:
3x-5*-3=21
3x=6
x=2
Solution:
x=2; y=-3
Cheers,
Stan H.